Solve the equation for r:
-3(6r+5)
______ = 6r-1
8


The line is a division ba

Answer by Abdul Hannan(9)

Solution ::
note :: the symbol * means multiplication and the symbol / means division
-3(6r+5)/8 = 6r-1
-3*(6r+5) = 8*(6r-1) [8 is transposed from the left side of the equal sign to its right side]
-3*6r - 3*5 = 8*6r - 8 [we now open the brackets]
-18r - 15 = 48r - 8
we now collect the like terms to the left side and right sides respectively
=> 48r+18r = -15+8
=> 66r = -7
=> r = -7/66